Question 1.39 Managertal accounting reports are 2 poiris related to the entire business enity orly prepared according to management needs O prepared according to GAAP O prepared periodically only. Question 1.40 rtod oosts ยอ $400,000 andthe unit contrtutonmwginissa. what amount or units must be sold in order to have a zero profit? O 10,000 unts O 400,000 units O 20,000 units Question 1.41 tod costs are S750,000 and variable coets are e0%of8ales, what 18the break4ven point saies dows? O $1,250,000 00.000 O $1.875 000 O $450.000 t Question 1.42 2 points O 2,000 units O 4808 units k Question 1.43 2 pairnts Richards Corporation had net income of $250,000 and paid dividende to common stockholdere of $sq,0o0. it had 50.000 shares of common stook outstanding uring the ebire year. Richards Corporations common slock te sealting for $35 per shar. The price earmings rato la: O 2imes
Answer #1

Answer 1.39 correct option (b)

Managerial accounting reports are prepared according to management needs

Answer 1.40 correct option (d) 20000 units

Explanation: Contribution margin is equal to Sales price per unit - Variable cost incurred per unit

Unit Contribution margin is given as $20 and fixed costs are $400000 given in question

we know at zero profit, Sales Revenue= Total Fixed costs + Total Variable costs

Total Fixed costs= Sales Revenue- Total Variable costs ( cost per unit * number of units)

Let the number of units produced be x, then $400000= 20x (at zero profit)

therefore solving for to get value of x i.e. $400000/$20 = 20000 units

Answer 1.41 correct option (c) 1875000

Calculation: At break even point (No profit, No loss situation), Sales Revenue = Fixed Costs + Variable Costs

Break even point in sales dollar is $1875000 because in this case Fixed + Variable costs are equal to Sales revenue

i.e. $750000 + 1125000( 60% of 1875000) = $1875000

Answer 1.42 correct option (d) 4808 units

Calculation: Let number of units be x

At break even point, Total fixed costs+ variable costs = Sales Revenue i.e. $250000+ $73x = 125x

$250000 = $52x

Solving for x, value of x = $ 250000/52 = 4807.69 round off i.e. 4808 units for break even

Answer 1.43 correct option (c) 7 times

Price-Earning Ratio (PE) is calculated by Price per share/ Earnings per share

Company's common stock is selling at $35 per share. Richard Corporation has 50000 outstanding shares and generated $250000 of net income. Hence Earning per share is equal to $250000/50000 = $5 per share

Hence Price Earning Ratio will be 7 times ($35 selling price per share/$5 earnings per share)